Evaluate the following.

1. 1^i

2.lm^ 2i+3

3.exp (i^3)

4.e^3iπ +1

5.(1+i)^10000


1
Expert's answer
2021-02-24T07:29:54-0500

1: Answer=1^i= 1

2:i^((2 * i) + 3) =

- 0.0432139183 i

3:. Answer=0.540302306 - 0.841470985 i

4:  e^^3iπ =-1 and e^1=2.71828182846 then the answer will be 2.71828182846X-1=-2.71828182846

5: 1^1000+i^1000 =1-1=0